OpenStudy (anonymous):

What is the exact volume of a sphere with a diameter of 18 cm? cm3 @iGreen I got 7776

OpenStudy (igreen):

No..I got something different. \(V = \dfrac{4}{3}r^3\) This is to find it in terms of pi..

OpenStudy (igreen):

Yeah

OpenStudy (igreen):

\(V = \dfrac{4}{3}(9^3)\) Simplify

OpenStudy (anonymous):

18x18x18=5832

OpenStudy (igreen):

You used 18..but that's the DIAMETER, and we need the radius, which is half the diameter. 18 / 2 = 9

OpenStudy (anonymous):

V≈3053.63cm³

OpenStudy (anonymous):

that's wrong it has to be an exact answer ....no decimals

OpenStudy (igreen):

Multiply 9 * 9 * 9 * 4/3

OpenStudy (anonymous):

ohh

OpenStudy (igreen):

That will be your answer.

OpenStudy (anonymous):

972

OpenStudy (anonymous):

May someone please help me?

OpenStudy (igreen):

Correct! \(972\pi\) is our answer :)

OpenStudy (anonymous):

@iGreen can you help me?